OpenStudy (anonymous):

how do you write the square root of -232 in terms of i

OpenStudy (compphysgeek):

with \(i^2 = -1\) you can say that \(-232 = i^2 \times 232\) and therefore \(\sqrt{-232} = i\,\sqrt{232}\)

OpenStudy (snuggielad):

See, i is the square root of a -1 In order to remove it you just take out the negative so it would be \[i \sqrt{232}\]

OpenStudy (snuggielad):

Then you just take the square root of 232 and write it beside the i leaving the remainder inside the radical

OpenStudy (snuggielad):

\[radical = \sqrt{(number here)} <---\]

OpenStudy (anonymous):

thank you

OpenStudy (snuggielad):

Don't forget to medal the best answer.